### v1

@@ IV-B1 The number of constraints
1 Input: lower bound $lb$ and upper bound vectors $ub$ where the length is prediction horizon T time step. Starting point $pt_{0}$ corresponding to the ego cars current position in the ST graph.
@@ IV-B2 Quick evaluation of constraints
As a pre-processing step we turn the bounds into monotonic functions, which is equivalent to making the assumption that our speed planner cannot go in reverse. These bounds are then used as inputs to the approximate profile from bounds algorithm depicted in Algorithm 1. The algorithm is a divide-and-conquer algorithm that splits the problem into sub-problems at the maximum violation of upper or lower bounds.
The algorithm starts by identifying the maximum safety margin, and then pads the bounds by that amount. The algorithm then calls the lower split algorithm described in Algorithm 2. Algorithm 2 and Algorithm 3 (in the appendix) are alternating checks that verify the problem or sub-problem satisfies the bounds. Algorithm 3 closely mirrors Algorithm 2, and was moved to the appendix for space. If the bounds are violated, the problem is decomposed into sub-problems where the bounds are guaranteed to be valid. Upper split and lower split are called recursively until all sub-problems are valid, at which point the sub-problems are returned and concatenated into the full solution. Because both upper split and lower split must be called at least once, there is a flag to ensure correctness. Proofs of the validity of the algorithm are included in Section -E of the Appendix.
@@ IV-B3 Pairing constraint sets
To reduce the combinatorial effect between the number of futures and the number of constraints we pair the constraint sets for each future. We pair constraints based on similarity of approximate trajectories. This is reasonable as largely diverse trajectories often produce infeasible or highly suboptimal solutions, however it is possible that the optimal solution is not the paired trajectory. The pairing heuristic reduces the $O(\bar{c}^{m})$ complexity to $O(\bar{c})$, however we lose our ability to guarantee optimality.
@@ IV-B4 Agent futures
Figure 4: Validation run on 1/10 scale RC cars. (Left) Shows the scenario, (Center) show an RVIZ visualization with predictions in yellow. (Right) Show the different plans corresponding to the different predictions.
Summarizing the impact of these approximations, we start with $\bar{c}^{m}=O(2^{k^{r^{k}}})$ problems of size $O(r^{k})$. Pairing constraint sets reduce this to $O(2^{k})$ problems of size $O(r^{k})$. In many real world driving scenarios, this is closer to $\Theta(k)$ problems of size $O(r^{k})$. The problem size is then reduced to $O(r^{l})$ where $l<4$ by restricting the number or agents with multiple futures. Some of these problems can be quickly discarded with a preliminary analysis made possible by Algorithm 1. We find that when $k<16$ and $m<8$ the system can run in realtime.
@@ V Validation
To validate our algorithm, we run our implementation in the simulator and on a 1/10 scale car. Traffic vehicles follow a constant velocity policy and GLK [isele2024gaussian] is used for prediction. Figure 2 shows an example of a scenario in CARLA. An MPC formulation that only predicts the car moving forward would be at risk if the car turns, or conversely, immediately slow down taking the possibility of a turn as given. The multi-future planner is able to handle the multiple predictions. We also observe, that because the decision is delayed, we do not immediately brake for a turning car, but are able to slow down enough so that we can brake in the future if needed. Figure 4 depicts a scenerio with 1/10 scale RC. In the plan, the agent waits long enough to determine if the car is passing, where it will then accelerate or continue to wait accordingly. An MPC formulation with only the (most likely) forward prediction will immediately accelerate and then not be able to respond when the prediction is updated.
@@
This section provides a visual example of a motivating problem for multi-future planning using a space-time (ST) graph. The left diagram in Fig. 5 shows an ST graph which allows us to visualize an agents temporal traversal along a path. Obstacles that block the path are depicted as rectangles, where the different obstacle correspond to two different possible futures. Feasible trajectories corresponding to the red obstacle are depicted in region A. Feasible trajectories corresponding to blue obstacle are depicted in region B. Given beliefs are fixed and we cannot influence other agents, there is no possible trajectory that will satisfy both possible futures. So in this case selecting the future that is more probable, and optimizing the trajectory in that corresponding region is optimal, but has an expected reward of $\infty$- ∞ since it will collide with a non-zero probability.
@@
The right diagram in Fig. 5 depicts a similar situation except now there is a safe trajectory that is more conservative, so if we were using the Expectation in Eqn. (3) we would choose the conservative trajectory from region C. However, since we know that only one future is possible, we could stay in the overlap of A and B, and make a decision at 4s, at which point we should know whether A or B corresponds to the true future.
@@
$\mathds{P}(f_{*}=f_{1})*R$ $(\tau_{t_{d}}|f_{1})+\mathds{P}(f_{*}=f_{2})*R(\tau_{t_{d}}|f_{2})\enspace.$ (15)
Where $\tau_{t_{d}}$ is that trajectory that results when making a decision at time $t_{d}$. Alternatively, if you make an earlier decision at alternate time $t_{a}$ (i.e., $t_{a}<t_{d}$), the expected reward is
@@
Case 2: we compare the case of the agent choosing $\tau_{2}$ at time $t_{d}$ vs. the case of the agent cohosing $\tau_{2}$ at time $t_{a}$ Here both agents choose a catastrophic trajectory
@@
Case 4: we compare the case of the agent choosing $\tau_{2}$ at time $t_{d}$ vs. the case of the agent choosing $\tau_{1}$ at time $t_{a}$. Given accuracy in belief is strictly non-decreasing (Assumption 4) and the agents are rational (i.e., a trajectory with higher belief is selected), this case is not possible.
In case 1, the agent has a potential loss of $k_{1}$, and in case 3 the agent has a potential gain of $k_{2}+\infty$. For an evaluation of the optimal reward, the likelihood of case 1 and case 3 would need to be determined, which will be scenario specific, but the analysis of the particular cases is sufficient to show that the possible benefit of an earlier decision is finite, where as the potential loss of an earlier decision is infinite. ∎
@@
Figure 6: QP problem growth. This contrived example that shows the worst-case scaling. Going above or below an obstacle results in $c=2^{k}$ possible quadratic problems for $k$ obstacles in the worst case.
@@ 17 $second\_half\leftarrow lower\_split(lb_{s\_id:end},ub_{s\_id:end},ub[s\_id],pt_{E},True)$
The constraints corresponding to the four futures are not convex. There can be $O(2^{k})$ constraints for each future. This upper bound is $322^{5}=$ 32 with five agents, however in practice this is often much lower based on the ordering of which agents cross our path and when. Figure 9 shows that in our example, there are three convex bounds $\bar{c}$ for each future.
@@
Figure 11: Making decisions based on probabilities at the first decision point can be greedy. For example, while future A is more probable, selecting future A possibly forces us to other choices which have comparatively lower likelihood.
@@
We believe that the probabilities of different futures can be conditioned on our action. and in this way we could expand our existing model to reason about probing actions for interactive decision making.
@@
In this work we do not consider the interactivity of driving. In driving, sometimes we deliberately make our intention clear. Selecting a trajectories that keeps the ego agent as flexible as possible can make it difficult for other agents to predict our action. We leave it as future work to explore how we can clear;y indicate our intention while remaining flexible.

### v2

@@ IV-B1 The number of constraints
1 Input: lower bound $lb$ and upper bound vectors $ub$ where the length is prediction horizon T time step. Starting point $pt_{0}$ corresponding to the ego car’s current position in the ST graph.
@@ IV-B2 Quick evaluation of constraints
As a pre-processing step we turn the bounds into monotonic functions, which is equivalent to making the assumption that our speed planner cannot go in reverse. These bounds are then used as inputs to the approximate profile from bounds in Algorithm 1. The algorithm is a divide-and-conquer algorithm that splits the problem into sub-problems at the maximum violation of upper or lower bounds.
The algorithm starts by identifying the maximum safety margin, and then pads the bounds by that amount. The algorithm then calls the lower split algorithm described in Algorithm 2. Algorithm 2 and Algorithm 3 (in the appendix) are alternating checks to verify if the problem or sub-problem satisfies the bounds. Algorithm 3 closely mirrors Algorithm 2, and was moved to the appendix for space. If the bounds are violated, the problem is decomposed into sub-problems where the bounds are guaranteed to be valid. Upper split and lower split are called recursively until all sub-problems are valid, at which point the sub-problems are returned and concatenated into the full solution. Because both the upper split and lower split must be called at least once, there is a flag to ensure correctness. Proofs of the validity of the algorithm are included in Section -E of the Appendix.
@@ IV-B3 Pairing constraint sets
To reduce the combinatorial effect between the number of futures and the number of constraints, we pair the constraint sets for each future. We pair constraints based on the similarity of approximate trajectories. This is reasonable as largely diverse trajectories often produce infeasible or highly suboptimal solutions, however, it is possible that the optimal solution is not the paired trajectory. The pairing heuristic reduces the $O(\bar{c}^{m})$ complexity to $O(\bar{c})$, but we lose our ability to guarantee optimality.
@@ IV-B4 Agent futures
Figure 4: Validation run on 1/10 scale RC cars. (Left) A snapshot of the scenario; (Center) An RVIZ visualization with predictions in yellow. (Right) A visualization of different plans corresponding to different predictions.
Summarizing the impact of these approximations, we start with $\bar{c}^{m}=O(2^{k^{r^{k}}})$ problems of size $O(r^{k})$. Pairing constraint sets reduce this to $O(2^{k})$ problems of size $O(r^{k})$. In many real world driving scenarios, this is closer to $\Theta(k)$ problems of size $O(r^{k})$. The problem size is then reduced to $O(r^{l})$ where $l<4$ by restricting the number or agents with multiple futures. Some of these problems can be quickly discarded with a preliminary analysis made possible by Algorithm 1. We find that when $k<16$ and $m<8$, the system can run in realtime.
@@ V Validation
To validate our algorithm, we run our implementation in the simulator and on a 1/10 scale car. Traffic vehicles follow a constant velocity policy and GLK [isele2024gaussian] is used for prediction. Figure 2 shows an example of a scenario in CARLA. An MPC formulation that only predicts the car moving forward would be at risk if the car turns, or conversely, immediately slows down, taking the possibility of a turn as given. The multi-future planner is able to handle multiple predictions. We also observe that because the decision is delayed, we do not immediately brake for a turning car, but are able to slow down enough so that we can brake in the future if needed. Figure 4 depicts a scenario with 1/10 scale RC. In the plan, the agent waits long enough to determine if the car is passing, where it will then accelerate or continue to wait accordingly. An MPC formulation with only the (most likely) forward prediction will immediately accelerate and then not be able to respond when the prediction is updated.
@@
This section provides a visual example of a motivating problem for multi-future planning using a space-time (ST) graph. The left diagram in Fig. 5 shows an ST graph which allows us to visualize an agents temporal traversal along a path. Obstacles that block the path are depicted as rectangles, where the different obstacle correspond to two different possible futures. Feasible trajectories corresponding to the red obstacle are depicted in region A. Feasible trajectories corresponding to the blue obstacle are depicted in region B. Given that beliefs are fixed and we cannot influence other agents, there is no possible trajectory that will satisfy both possible futures. So in this case, selecting the more probable future and optimizing the trajectory in that corresponding region is optimal, but has an expected reward of $\infty$- ∞ since it will collide with a non-zero probability.
@@
The right diagram in Fig. 5 depicts a similar situation, except now there is a safe trajectory that is more conservative, so if we were using the Expectation in Eqn. (3), we would choose the conservative trajectory from region C. However, since we know that only one future is possible, we could stay in the overlap of A and B, and make a decision at 4s, at which point we should know whether A or B corresponds to the true future.
@@
$\mathds{P}(f_{*}=f_{1})*R$ $(\tau_{t_{d}}|f_{1})+\mathds{P}(f_{*}=f_{2})*R(\tau_{t_{d}}|f_{2})\enspace,$ (15)
where $\tau_{t_{d}}$ is the trajectory that results when making a decision at time $t_{d}$. Alternatively, if you make an earlier decision at alternate time $t_{a}$ (i.e., $t_{a}<t_{d}$), the expected reward is
@@
Case 2: we compare the case of the agent choosing $\tau_{2}$ at time $t_{d}$ vs. the case of the agent cohosing $\tau_{2}$ at time $t_{a}$ Here, both agents choose a catastrophic trajectory
@@
Case 4: we compare the case of the agent choosing $\tau_{2}$ at time $t_{d}$ vs. the case of the agent choosing $\tau_{1}$ at time $t_{a}$. Given that accuracy in belief is strictly non-decreasing (Assumption 3) and the agents are rational (i.e., a trajectory with higher belief is selected), this case is not possible.
In case 1, the agent has a potential loss of $k_{1}$, and in case 3, the agent has a potential gain of $k_{2}+\infty$. For an evaluation of the optimal reward, the likelihood of case 1 and case 3 would need to be determined, which will be scenario specific, but the analysis of the particular cases is sufficient to show that the possible benefit of an earlier decision is finite, where as the potential loss of an earlier decision is infinite. ∎
@@
Figure 6: QP problem growth. This contrived example shows the worst-case scaling. Going above or below an obstacle results in $c=2^{k}$ possible quadratic problems for $k$ obstacles in the worst case.
@@ 17 $second\_half\leftarrow lower\_split(lb_{s\_id:end},ub_{s\_id:end},ub[s\_id],pt_{E},True)$
The constraints corresponding to the four futures are not convex. There can be $O(2^{k})$ constraints for each future. This upper bound is $322^{5}=$ 32 with five agents, however in practice, this is often much lower based on the ordering of which agents cross our path and when. Figure 9 shows that in our example, there are three convex bounds $\bar{c}$ for each future.
@@
Figure 11: Making decisions based on probabilities at the first decision point can be greedy. For example, while future A is more probable, selecting future A possibly forces us to other choices that have a comparatively lower likelihood.
@@
We believe that the probabilities of different futures can be conditioned on our action. In this way, we could expand our existing model to reason about probing actions for interactive decision making.
@@
In this work, we do not consider the interactivity of driving. In driving, sometimes we deliberately make our intention clear. Selecting a trajectory that keeps the ego agent as flexible as possible can make it difficult for other agents to predict our action. We leave it as future work to explore how we can clear;y indicate our intention while remaining flexible.